What is the x-intercept of the line with the equation 4x+2y=12?: What is the x-intercept of the line with the equation 4x+2y=12?
Answer by Earlsdon(3748)

Find the x-intercept of:
4x+2y = 12
Remember, the x-intercept occurs at y = 0, so set y = 0 and solve for x.
4x+2(0) = 12
4x+0 = 12
4x = 12 Divide both sides by 4.
x = 12/4
x = 3 This is the x-intercept!

Check:
4x+2y = 12
4(3)+2y = 12
12 + 2y = 12 Subtract 12 from both sides.
2y = 0
y = 0
Check the graph of the equation;
graph(300,200,-5,5,-5,8,-2x+6)